# Enhancing Underwater Imaging with 4-D Light Fields: Dataset and Method

### v1

@@ Abstract
In this paper, we delve into the realm of 4-D light fields (LFs) to enhance underwater imaging plagued by light absorption, scattering, and other challenges. Contrasting with conventional 2-D RGB imaging, 4-D LF imaging excels in capturing scenes from multiple perspectives, thereby indirectly embedding geometric information. This intrinsic property is anticipated to effectively address the challenges associated with underwater imaging. By leveraging both explicit and implicit depth cues present in 4-D LF images, we propose a progressive, mutually reinforcing framework for underwater 4-D LF image enhancement and depth estimation. Specifically, our framework explicitly utilizes estimated depth information alongside implicit depth-related dynamic convolutional kernels to modulate output features. The entire framework decomposes this complex task, iteratively optimizing the enhanced image and depth information to progressively achieve optimal enhancement results. More importantly, we construct the first 4-D LF-based underwater image dataset for quantitative evaluation and supervised training of learning-based methods, comprising 75 underwater scenes and 3675 high-resolution 2K pairs. To craft vibrant and varied underwater scenes, we build underwater environments with various objects and adopt several types of degradation. Through extensive experimentation, we showcase the potential and superiority of 4-D LF-based underwater imaging vis-a-vis traditional 2-D RGB-based approaches. Moreover, our method effectively corrects color bias and achieves state-of-the-art performance. The dataset and code will be publicly available at https://github.com/linlos1234/LFUIE.
@@ I Introduction
Over the past years, many physical model-based and learning-based enhancement methods [2, 3, 4, 5, 6, 7] have been proposed for improving the quality of underwater images obtained by traditional 2-D RGB imaging. These physical model-based approaches [2, 3, 4, 5] mainly concentrate on the depth-related parameters to the degradation process of the underwater imaging, while the deep learning algorithms [6, 7] apply the popular convolutional or transformer network architecture to overfit 2-D RGB-based dataset. However, these traditional 2-D RGB-based underwater imaging approaches still face many challenges, including complex and ill-posed optimization for model-based methods and the lack of realistic pairwise data for learning-based methods. Moreover, monocular depth estimations from 2-D RGB imaging are often inaccurate, limiting the performance of these methods.
@@
4-D light field (LF) imaging is an advanced multi-dimensional technique [8, 9] that can capture the same scene from multiple perspectives in a single shot. This technology can simultaneously record both the reflected light intensity and rich geometric information of the scene. Under the assumption of Lambertian, projections of the same scene point have the same intensity across different views. This geometric relation leads to a particular LF parallax structure, which makes depth estimation easier compared to traditional 2-D RGB imaging. In addition to this, LF images contain implicit geometric information, which allows for the aggregation of effective information from other dimensions. By aligning features, images from different perspectives can merge light rays and reduce noise while preserving geometric structures, resulting in improved consistency. Therefore, it is promising to explore the acquisition of high-quality underwater images through 4-D LF imaging.
@@
4-D LF images captured in underwater environments are essential for exploring this potential direction. Unfortunately, there is currently no publicly available dataset for such imagery. While LF cameras could be employed for data acquisition, this approach is laden with challenges, such as high expenses and the absence of authentic ground truth data. This scarcity complicates the quantitative assessment of LF technology’s potential in underwater imaging and also hinders supervised learning of networks. Consequently, we opt to synthesize lifelike underwater 4-D LF images through 3-D modeling and rendering software. Specifically, we utilize Blender [10] to conduct the first 4-D LF-based underwater image dataset that includes LF images, ground truth images, and depth maps, as depicted in Fig. 2. Furthermore, we propose a learning-based method to enhance underwater 4-D LF images. Our key innovation lies in that depth estimation and enhancement results promote each other. Our method enhances 4-D LF images by iterating multiple stages with both explicit and implicit geometric information. Specifically, we utilize the estimated depth maps from 4-D LF images to modulate the features. The geometric information extracted from epipolar plane image (EPI) features can also help adjust the spatial and angular features. Extensive experiments and ablative studies demonstrate the advantages of our approach.
@@
  • we build the first 4-D LF-based underwater image dataset LFUB, which contains 75 sets of realistic and high-resolution underwater 4-D LF images with corresponding ground truth no-water LF images and depth maps. Our LFUB dataset is expected to advance research in the field of underwater vision, and the pipeline of conducting is feasibly used to build other helpful LF-based datasets;
@@ II-A Underwater 2-D RGB Image/Video Enhancement
2) Physical Model-based Methods. Widely used physical model-based methods treat underwater image enhancement as an inverse problem, which utilizes prior knowledge to estimate the parameters of underwater imaging models. Researchers typically improve underwater images through the following sequential steps: defining the initial physical imaging model, estimating crucial parameters, and solving the inversion problem. The foundation of these methods lies in designing appropriate priors based on a physical understanding of the underwater world to accurately characterize the environment. These priors involve dark channel prior [17], red channel prior [18], minimum information prior [3, 4], haze-line prior [5] and etc.
Carlevaris-Bianco et al. [19] noted that red light often attenuates faster than green and blue light in underwater environments. Based on this observation, they designed a prior that leverages this differential attenuation to estimate transmission maps of underwater images. Drews et al. [20] proposed the Underwater Dark Channel Prior (UDCP) to address the selective attenuation of light. Li et al. [3] enhanced underwater images by employing a dehazing technique that minimizes information loss, incorporating a histogram distribution prior to improve image details. In addition to the minimum information prior, Li et al. [4] considered the specific optical properties of underwater environments. Berman et al. [5] adapted the haze-line prior for underwater images, which is based on the concept that different wavelengths are absorbed at different rates, causing colors to shift primarily towards blue and green. Moreover, Akkaynak and Treibitz [21] revised the underwater image formation model [22] and presented Sea-Thru, a physically accurate color correction method.
@@
Initially, due to the scarcity of pairwise data, Li et al. [26] employed WaterGAN and an imaging formation model to synthesize pairwise data with clean images and depth maps in an unsupervised manner. Subsequently, they adopted a two-stage training strategy to restore underwater images, especially for color correction. Li et al. [27] proposed a weakly supervised method, Water CycleGAN based on Cycle Consistent Adversarial Networks [28] and a multi-term loss function, relaxing the need for pairwise training data of underwater images. More recently, Li et al. [6] introduced the first pairwise underwater image dataset, which greatly advanced the performance of supervised learning and proposed WaterNet as a baseline network. Liu et al. [7] designed an adaptive learning module to learn representative features and transfer shallow information by residual learning. Uncertain learning [29] was utilized to help improve enhancement performance. Guo et al. [30] pre-trained a ranking-based model using a histogram prior and cross-scale correspondence to determine which input is better, then trained a simple enhancement network with pre-trained rank loss to achieve optimal outputs. Peng et al. [31] used a UNet shape transformer network with channel-wise and spatial-wise feature modulation and proposed a multi-color space loss function to improve image contrast. Moreover, Li et al. [32] used physical transmission maps as guidance to modulate features and exploited information from multiple color spaces to improve image quality.
In addition to single image methods, there are also some approaches for video-based underwater enhancement [33, 34], which leverages additional temporal information for improved results. Xie et al. [34] exploited the inter-frame information to optimize the enhancement guided by the information from the current frame.
@@ III-A Scene Construction
While adjusting the visual properties of the water, we make extensive use of Blender’s advanced shader capabilities, particularly focusing on the Principled Volume shader, as shown in Fig. 3. By manipulating the Anisotropy parameter, we are able to control the direction bias of specular reflections, allowing highlights to stretch along predominant directions. This adjustment closely mimics the way light behaves as it travels through water, which is often non-uniform and directionally biased due to factors like currents and particulate concentration. Additionally, we modify the Density parameter to alter opacity and the degree of light scattering within the water volume. Areas with higher density settings became more opaque, accurately reflecting the diverse visibility conditions commonly encountered in underwater environments.
@@ III-B Camera Setting
(a) Types of color deviation degradation for two datasets
@@
Figure 4: Statistics of the constructed underwater LF dataset. (a) Types of color deviation degradation for two datasets: the left one is RGB enhanced dataset UVBE [34], and the right one is our LFUB dataset. (b) Scene object categories.
@@ III-C Image Rendering
Figure 5: The variation in the degree of degradation aligns with human perception, and underwater degradation indeed diminishes as depth approaches, consistent with physical processes in Blender. The non-reference metrics UIQM and UCIQE increase as the visual quality improves. The first row shows that as the camera continuously moves forward, the degree of underwater degradation weakens for the same scene, and correspondingly the metrics become better. The second row shows that the metrics with the same color deviation increase as the degree of color deviation decreases.
@@ III-D Data Format
Figure 6: Two different forms of 4-D LF representation for one scene Plane. (a) Sub-aperture images: the content of each magnified area varies with the view position; (b) Epipolar plane images: slices of 4-D LF.
@@ III-E Statistics of LFUB Dataset
Figure 7: Network architecture of our multi-stage network. In each stage, we estimate the central view disparity from the input of this stage with OccUnNet [54]. ’DownSampling’ is implemented by max pooling, while ’UpSampling’ is implemented by average pooling.
@@ IV Proposed Method
In this section, we propose a learning-based method for enhancing underwater 4-D LF images, as illustrated in Fig. 7. The framework follows the basic idea that depth estimation and enhancement process depend on each other, and the network enhances underwater 4-D LF images through iterative stages. Specifically, in the initial stage, we feed the input to a multi-pattern feature interaction module (MFIM) at different scales. We then concatenate the multi-scale features passing to the depth module with the disparity from an estimator. In the depth modulation block, the former features are adjusted using explicit depth guidance and implicit EPI geometry convolution kernels. After modulation, the output features are passed through a convolution layer to reconstruct a result. In subsequent stages, the coarse result from the preceding stage is used as input, and the intermediate features from the previous stage are used for skip connection.
@@ IV-C Depth Modulation
According to Eq. (3), it is clear that depth plays a crucial role in influencing the extent of degradation across various regions. Thus, we explore both explicit modulation (EM) and EPI-guided implicit modulation (EPI-IM) of geometry information for 4-D LF data, as illustrated in Fig. 9. Explicitly, we integrate depth into the network by passing the disparity of the central view through a 3 $\times$ 3 convolution layer, generating a deformation map $\boldsymbol{\theta}$ to modulate the features. Rather than utilizing depth solely for guiding feature extraction, we employ depth to direct the restoration at the end of each stage, aligning with the underwater degradation. The explicit modulation process can be formulated as
@@
Figure 9: The schematic illustration of the Depth Modulation module. Each depth modulation module is composed of two MFIM blocks, one EPI-IM, and one EM.
Furthermore, recognizing that EPIs indirectly encode geometric information, we create convolutional kernels (EPIK) utilizing EPI features. Subsequently, spatial and angular features are adjusted based on geometric structural information, as depicted in Fig. 8. The implicit adjustment process is formulated as
@@
TABLE II: Quantitative comparisons (PSNR/SSIM/inference time) of different methods on proposed LFUB dataset. The best second best results are highlighted in bold and underline. “↑” (resp. “↓”) means the larger (resp. smaller), the better. All compared methods are retrained on our training set and evaluated on the testing set
    Methods    Fusion [12]    GDCP [60]    MMLE [61]    WWPF [62]    LANet [7]    PUIE [29]
[…]
Inf. $\downarrow$ 0.3512 0.6469 0.2894 0.7557 0.1122 0.2823
    Methods    Ushape [31]    Uranker [30]    UVE [33]    DistgSSR [53]    MSPNet [38]    Ours
[…]
Inf. $\downarrow$ 0.1836 0.1727 0.2449 1.0269 0.9196 0.6987
@@ V Experiments
V-A Implementation Details
We found that disparity estimation with the constructed 4-D LF underwater data is challenging due to the rugged nature of the scenes in our LFUB dataset. To address this, we pre-trained OccUnNet on a synthetic LF dataset, HCI Dataset [65], and then performed a warm-up operation on our 4-D LF underwater data by warping the side views to the central view.
V-B Experiment Settings

### v2

@@ Abstract
In this paper, we delve into the realm of 4-D light fields (LFs) to enhance underwater imaging plagued by light absorption, scattering, and other challenges. Contrasting with conventional 2-D RGB imaging, 4-D LF imaging excels in capturing scenes from multiple perspectives, thereby indirectly embedding geometric information. This intrinsic property is anticipated to effectively address the challenges associated with underwater imaging. By leveraging both explicit and implicit depth cues present in 4-D LF images, we propose a progressive, mutually reinforcing framework for underwater 4-D LF image enhancement and depth estimation. The entire framework decomposes this complex task, iteratively optimizing the enhanced image and depth information to progressively achieve optimal enhancement results. More importantly, we construct the first 4-D LF-based underwater image dataset for quantitative evaluation and supervised training of learning-based methods, comprising 75 underwater scenes with multiple views and 3675 high-resolution 2K pairs. To craft vibrant and varied underwater scenes, we build underwater environments with various objects and adopt several types of degradation. Through extensive experimentation, we showcase the potential and superiority of 4-D LF-based underwater imaging vis-a-vis traditional 2-D RGB-based approaches. Moreover, our method effectively corrects color bias and achieves state-of-the-art performance. The dataset and code will be publicly available at https://github.com/linlos1234/LFUIE.
@@ I Introduction
Over the past years, many physical model-based and learning-based enhancement methods [52, 42, 43, 44, 19, 29] have been proposed for improving the quality of underwater images obtained by traditional 2-D RGB imaging. These physical model-based approaches [52, 42, 43, 44] mainly concentrate on the depth-related parameters to the degradation process of the underwater imaging, while the deep learning algorithms [19, 29] apply the popular convolutional or transformer network architecture to overfit 2-D RGB-based datasets. However, these traditional 2-D RGB-based underwater imaging approaches still face many challenges, including complex and ill-posed optimization for model-based methods and the lack of realistic pairwise data for learning-based methods. Moreover, monocular depth estimations from 2-D RGB imaging are often inaccurate, causing limited performance.
@@
4-D light field (LF) imaging is an advanced multi-dimensional technique [8, 9] that can capture the same scene from multiple perspectives in a single shot. This technology can simultaneously record both the reflected light intensity and rich geometric information of the scene. Under the assumption of Lambertian, projections of the same scene point have the same intensity across different views. This geometric relation leads to a particular LF parallax structure, which makes depth estimation easier compared to traditional 2-D RGB imaging. In addition, LF images contain implicit geometric information, enabling the aggregation of valuable content from all views. By aligning features, images from different perspectives can merge light rays and reduce noise while preserving geometric structures, resulting in improved consistency. Therefore, it is promising to explore the acquisition of high-quality underwater images through 4-D LF imaging.
@@
4-D LF images captured in underwater environments are essential for exploring this potential direction. Unfortunately, there is currently no publicly available dataset for such imagery. While LF cameras could be employed for data acquisition, this approach is laden with challenges, such as high expenses and the absence of authentic ground truth data. This scarcity complicates the quantitative assessment of LF technology’s potential in underwater imaging and also hinders supervised learning of networks. Consequently, we opt to synthesize lifelike underwater 4-D LF images through 3-D modeling and rendering software. Specifically, we utilize Blender [72] to conduct the first 4-D LF-based underwater image dataset that includes LF images, ground truth images, and depth maps, as depicted in Fig. 2. Furthermore, we propose a learning-based method to enhance underwater 4-D LF images. Our key innovation lies in that depth estimation and enhancement results promote each other. The method enhances 4-D LF images by iterating multiple stages with both explicit and implicit geometric information. Specifically, we utilize the estimated depth maps from 4-D LF images to modulate the features. The geometric information extracted from epipolar plane image (EPI) features can also help adjust the spatial and angular features. Extensive experiments and ablative studies demonstrate the advantages of our approach.
@@
  • we build the first 4-D LF-based underwater image dataset LFUB, which contains 75 multi-view scenes of realistic and high-resolution underwater 4-D LF images with corresponding no-water LF images and depth maps. Our LFUB dataset is expected to advance research in the field of underwater vision, and the pipeline of conducting is feasibly used to build other helpful LF-based datasets;
@@ II-A Underwater 2-D RGB Image/Video Enhancement
2) Physical Model-based Methods. Widely used physical model-based methods treat underwater image enhancement as an inverse problem, which utilizes prior knowledge to estimate the parameters of underwater imaging models. Researchers typically improve underwater images through the following sequential steps: defining the initial physical imaging model, estimating crucial parameters, and solving the inversion problem. The foundation of these methods lies in designing appropriate priors based on a physical understanding of the underwater world to accurately characterize the environment. These priors involve dark channel prior [39], red channel prior [41], minimum information prior [42, 43], haze-line prior [44], etc.
Carlevaris-Bianco et al. [38] noted that red light often attenuates faster than green and blue light in underwater environments. Based on this observation, they designed a prior that leverages this differential attenuation to estimate transmission maps of underwater images. Drews et al. [40] proposed the Underwater Dark Channel Prior (UDCP) to address the selective attenuation of light. Li et al. [42] enhanced underwater images by employing a dehazing technique that minimizes information loss, incorporating a histogram distribution prior to improve image details. In addition to the minimum information prior, Li et al. [43] considered the specific optical properties of underwater environments. Berman et al. [44] adapted the haze-line prior for underwater images, which is based on the concept that different wavelengths are absorbed at different rates, causing colors to shift primarily towards blue and green. Zhuang et. al [74] proposed a retinex variational model using hyper-Laplacian reflectance priors, which applied sparsity-promoting and complete-comprehensive reflectance. Moreover, Akkaynak and Treibitz [21] revised the underwater image formation model [2] and presented Sea-Thru, a physically accurate color correction method.
@@
Initially, due to the scarcity of pairwise data, Li et al. [45] employed WaterGAN and an imaging formation model to synthesize pairwise data with clean images and depth maps in an unsupervised manner. Subsequently, they adopted a two-stage training strategy to restore underwater images, especially for color correction. Li et al. [46] proposed a weakly supervised method, Water CycleGAN based on Cycle Consistent Adversarial Networks [47] and a multi-term loss function, relaxing the need for pairwise training data of underwater images. More recently, Li et al. [19] introduced the first pairwise underwater image dataset, which greatly advanced the performance of supervised learning and proposed WaterNet as a baseline network. Liu et al. [29] designed an adaptive learning module to learn representative features and transfer shallow information by residual learning. Uncertain learning [28] was utilized to help improve enhancement performance. Guo et al. [3] pre-trained a ranking-based model using a histogram prior and cross-scale correspondence to determine which input is better, then trained a simple enhancement network with pre-trained rank loss to achieve optimal outputs. Peng et al. [4] used a UNet shape transformer network with channel-wise and spatial-wise feature modulation and proposed a multi-color space loss function to improve image contrast. Wang et al. [75, 76] enhanced the underwater image quality by mining internal and external information and using domain adaptation and representation learning methods. Moreover, Li et al. [51] used physical transmission maps as guidance to modulate features and exploited information from multiple color spaces to improve image quality.
In addition to single image methods, there are also some approaches for video-based underwater enhancement [5, 23], which leverage additional temporal information for improved results. Xie et al. [23] exploited the inter-frame information to optimize the enhancement guided by the information from the current frame.
@@ III-A Scene Construction
While adjusting the visual properties of the water, we make extensive use of Blender’s advanced shader capabilities, particularly focusing on the Principled Volume shader, as shown in Fig. 3. By manipulating the Anisotropy parameter, we can control the direction bias of specular reflections, allowing highlights to stretch along predominant directions. This adjustment closely mimics the way light behaves as it travels through water, which is often non-uniform and directionally biased due to factors like currents and particulate concentration. Additionally, we modify the Density parameter to alter opacity and the degree of light scattering within the water volume. Areas with higher density settings became more opaque, accurately reflecting the diverse visibility conditions commonly encountered in underwater environments. We show two groups of images with different Anisotropy and Density, in Fig. 4. Specifically, we change the Density from 0.02 to 0.17 in the first line with the fixed Anisotropy. And we change the Anisotropy from 0 to 1.2 in the second line with the fixed Density. As we can see, the Density controls the lightness of the water and the Anisotropy controls the color of the water.
@@
(a) D=0.02
(b) D=0.07
(c) D=0.12
(d) D=0.17
(e) A=0.0
(f) A=0.4
(g) A=0.8
(h) A=1.2
Figure 4: Visual comparisons of different Anisotropy (A) and Density (D). Zoom in for details.
@@ III-B Camera Setting
(a) Types of color deviation degradation for two datasets. Left: 2-D RGB dataset UVBE [23]. Right: 4-D LF dataset LFUB
@@
Figure 5: Statistics of the constructed underwater LF dataset. (a) Types of color deviation degradation for two datasets: the left one is 2-D RGB enhanced dataset UVBE [23], and the right one is our 4-D LFUB dataset. (b) Scene object categories.
@@ III-D Data Format
Figure 6: The variation in the degree of degradation aligns with human perception, and underwater degradation indeed diminishes as depth approaches, consistent with physical processes in Blender. The non-reference metrics UIQM and UCIQE increase as the visual quality improves. Zoom in for details.
Figure 7: Two different forms of 4-D LF representation for one scene Plane. (a) Sub-aperture images: the content of each magnified area varies with the view position; (b) Epipolar plane images: slices of 4-D LF. Zoom in for details.
@@ III-E Statistics of LFUB Dataset
Figure 8: Network architecture of our multi-stage network. In each stage, we estimate the central view disparity from the input of this stage with OccUnNet [27]. In the enhancement branch, we first extract features by multi-pattern convolution at different scales. Then, the multi-pattern features are passed to the Depth Modulation Block. We estimate dynamic kernels from EPI features, which are used to adjust spatial and angular features. At Last, the features are enhanced by the estimated disparities. The network then outputs the results from the enhanced features.
@@ IV Proposed Method
In this section, we propose a learning-based method for enhancing underwater 4-D LF images, as illustrated in Fig. 8. The framework follows the basic idea that depth estimation and the enhancement process depend on each other, and the network enhances underwater 4-D LF images through iterative stages. Specifically, in the initial stage, we feed the input to a multi-pattern feature interaction module (MFIM) at different scales. We then concatenate the multi-scale features passing to the depth module with the disparity from an estimator. In the depth modulation block, the former features are adjusted using explicit depth guidance and implicit EPI geometry convolution kernels. After modulation, the output features are passed through a convolution layer to reconstruct a result. In subsequent stages, the coarse result from the preceding stage is used as input, and the intermediate features from the previous stage are used for skip connection.
@@ IV-C Depth Modulation
According to Eq. (3), it is clear that depth plays a crucial role in influencing the extent of degradation across various regions. The 4D structure is equivalent to implicitly recording the depth information. Thus, we explore both explicit modulation (EM) and EPI-guided implicit modulation (EPI-IM) of geometry information for 4-D LF data, as illustrated in Fig. 10. Explicitly, we integrate depth into the network by passing the disparity of the central view through a 3 $\times$ 3 convolution layer, generating a deformation map $\boldsymbol{\theta}$ to modulate the features. Rather than utilizing depth solely for guiding feature extraction, we employ depth to direct the restoration at the end of each stage, aligning with the underwater degradation. The explicit modulation process can be formulated as
@@
Furthermore, recognizing that EPIs indirectly encode geometric information [77], we create convolutional kernels (EPIK) utilizing EPI features. Subsequently, spatial and angular features are adjusted based on geometric structural information, as depicted in Fig. 9. The implicit adjustment process is formulated as
@@ IV-D Progressive Enhancement and Disparity Estimation
Figure 10: The schematic illustration of the Depth Modulation module. Each depth modulation module is composed of two MFIM blocks, one EPI-IM, and one EM. DM block first estimates multi-pattern features and then generates dynamic kernels from EPI features to adjust spatial and angular features. At Last, the features are enhanced based on the estimated disparities.
@@ IV-E Loss Function
TABLE II: Quantitative comparisons (PSNR/SSIM/inference time) of different methods on proposed LFUB dataset. The best second-best results are highlighted in bold and underline. “↑” (resp. “↓”) means the larger (resp. smaller), the better. All compared methods are retrained on our training set and evaluated on the testing set
    Methods    Fusion [33]    GDCP [60]    MMLE [58]    WWPF [59]    LANet [29]    PUIE [28]
[…]
Inf. $\downarrow$ 0.3512 0.6469 0.2894 0.7557 0.1122 0.2823
    Methods    Ushape [4]    Uranker [3]    UVE [5]    DistgSSR [30]    MSPNet [6]    Ours
[…]
Inf. $\downarrow$ 0.1836 0.1727 0.2449 1.0269 0.9196 0.6987
@@ V Experiments
V-A Implementation Details
We found that disparity estimation with the constructed 4-D LF underwater data is challenging due to the rugged nature of the scenes in our LFUB dataset. To address this, we pre-trained OccUnNet on a synthetic LF dataset, HCI Dataset [57], and then performed a warm-up operation on our 4-D LF underwater data by warping the side views to the central view. We chose this network because it leverages LF parallax to deliver precise relative depth estimation, while its compact size enables efficient joint training with the main network.
V-B Experiment Settings